If your investment doubles in 6​ 3/4 years, what approximate annual rate of return would you have​ earned? If you could earn an annual rate of​ 7.50%, approximately how long would it take for your investment to​ double?

A. ​9.60%; 10.67 years

B. ​13.50%; 9.20 years

C. ​9.20%; 13.50 years

D. ​10.67%; 9.60 years

answer is NOT 10.81 and 9.58 !
Answer #1
A B C D E F G H I J
2
3 a)
4
5 Let n be the number of years amount P is invested at rate APR at compounding frequency f, then
6 Future value (FV) can be calculated as follows:
7 FV = P*(1+ APR/f)n
8
9 In case of investment is doubled then
10 FV = 2P
11
12 Then,
13 2P = P*(1+ APR/f)n
14 or
15 2 = (1+ APR/f)n
16 or
17 APR =(21/n-1)*f
18
19 Given
20 f 1 (Annual Compounding)
21 n 6.75
22
23 APR =(21/n-1)*f
24 10.81% =((2^(1/D21))-1)*D20
25
26 Hence APR is 10.81%
27
28
29 Calculation of time required to double the investment:
30
31 As derived above equation for doubling of investment is
32 2 = (1+ APR/f)n
33 or
34 n =ln(2) / ln(1+APR/f)
35
36 Given
37 f 1 (Annual Compounding)
APR 7.50%
39
40 n =ln(2) / ln(1+APR/f)
41 9.58 =LN(2) / LN(1+D38/D37)
42
43 Hence Number of Years is 9.58
44
45 Thus the option (D) is approximately correct
46